OpenStudy (anonymous):

a grid shows the positions of a subway stop and your house. the subway stop is located at (7, -6) and your house is located at (-9, 5) what is the distance, to the nearest unit, betweem your house and the subway stop?

OpenStudy (dumbcow):

Think of it as a triangle and the distance is the hypotenuse the horizontal leg of triangle has length(7-(-9)) or 16 vertical leg of triangle has length (5-(-6)) or 11 use Pythagorean theorem c^2 = 16^2 + 11^2 c^2 = 256 + 121 c^2 = 377 c = sqrt(377) = 19.4